# Linear-Time Demonstration Selection for In-Context Learning via Gradient Estimation

### v1

@@ Abstract
This paper introduces an algorithm to select demonstration examples for in-context learning of a query set. Given a set of $n$ examples, how can we quickly select $k$ out of $n$ to best serve as the conditioning for downstream inference? This problem has broad applications in prompt tuning and chain-of-thought reasoning. Since model weights remain fixed during in-context learning, previous work has sought to design methods based on the similarity of token embeddings. This work proposes a new approach based on gradients of the output taken in the input embedding space. Our approach estimates model outputs through a first-order approximation using the gradients. Then, we apply this estimation to multiple randomly sampled subsets. Finally, we aggregate the sampled subset outcomes to form an influence score for each demonstration, and select $k$ most relevant examples. This procedure only requires pre-computing model outputs and gradients once, resulting in a linear-time algorithm relative to model and training set sizes. Extensive experiments across various models and datasets validate the efficiency of our approach. We show that the gradient estimation procedure yields approximations of full inference with less than $\mathbf{1}\%$ error across six datasets. This allows us to scale up subset selection that would otherwise run full inference by up to $\mathbf{37.7}\times$ on models with up to $34$ billion parameters, and outperform existing selection methods based on input embeddings by $\mathbf{11}\%$ on average.
@@
footnotetext: *Equal Contribution. Email correspondence can be directed to all authors at {zhang.zini, zhang.zhens, li.dongyu, ho.zhang}@northeastern.edu, wangluxy@umich.edu, and jdy@ece.neu.edu.footnotetext: *Equal Contribution. Email correspondence can be directed to all authors at {zhang.zini, zhang.zhens, li.dongyu, ho.zhang}@northeastern.edu, wangluxy@umich.edu, and jdy@ece.neu.edu.
@@ 1 Introduction
Table 1: Runtime comparison between our algorithms and standard subset selection methods for demonstration selection. Here, $T$ represents the time cost of running one forward pass on a model with the training set, $k$ denotes the length of the selected prompts, and $n$ represents the number of demonstrations. Our approach computes model outputs with one forward pass of time $T$ and input gradients with time $2T$, plus an overhead cost of $O(n\log n)$, which is negligible relative to $T$.
@@ 2 Preliminaries
Our goal is to select a subset $S\subseteq\mathcal{S}^{\text{Demo}}$ that minimizes the above objective. After selecting a set of demonstrations, we evaluate the test performance on a separate test set.
@@ 3 Scaling Up Inference via Estimation
A case study of linear functions. We illustrate the above approximation with an example of learning linear functions with transformer models Garg et al. (2022). We use this setting to evaluate the accuracy of equation (1) for ICL of linear functions. Suppose that both the in-context example and query example are generated as $y_{i}=\left\langle\beta^{(j)},x_{i}\right\rangle+\epsilon_{i}$, where the input vector $x_{i}$ is randomly sampled from $\mathcal{N}(0,\operatorname{Id}_{d})$ with $d=20$. The coefficient vector $\beta^{(j)}$ is randomly drawn from a finite set $\{\beta^{(1)},\beta^{(2)},\cdots,\beta^{(C)}\}$ with $C<d$, meaning that different in-context examples may have different coefficients. We consider two settings for the noise term $\epsilon_{i}$: linear regression with $\epsilon_{i}=0$ for all $i$, or noisy linear regression with $\epsilon_{i}\sim\mathcal{N}(0,1)$.
Figure 2(a) shows that in both settings, ICL-GradEx can accurately estimate the true outcomes of ICL. In particular, as $k$ increases, the trained transformer learns the linear function correctly, and the residual sum of squares (RSS) between the estimated scores and the true scores gradually decreases to near zero. Recall that $d=20$. We thus plot from $k=d$ onward because this is the minimum number of samples required to solve a system of linear equations. When $k<d$, the trained transformer is unable to fully determine the linear function. (See Figure 2, Garg et al. (2022)).
@@ 4 Efficient Demonstration Selection
Figure 2: An illustration of our approach for in-context learning of linear functions, as we vary the number of in-context examples $k$. Figure 2(a): ICL-GradEx incurs low approximation error relative to full inference, for both linear regression and noisy linear regression. Figure 2(b): With ICL-GradSel, the selected demonstrations follow the same linear function $\beta$, achieving lower error than top-$k$ and random-$k$ selections.
@@ 5 Experiments
Our experiments show that, across a wide range of LLMs (such as Llama, OPT, DeepSeek, and Qwen), ICL-GradEx can estimate model inference outcomes within $\textbf{1}\%$ error, for LLMs ranging from $1$ billion up to $34$ billion parameters. Crucially, this estimation allows us to reduce the computation cost of subset selection by up to $\textbf{37.7}\times$. Applied to in-context learning evaluations across six datasets that include both sentence classification and math reasoning tasks, ICL-GradRE improves the test $F_{1}$ score by $\textbf{11}\%$ on average. For long-context settings where $k=150$, our selection matches the performance of top-$k$ and random-$k$ selection with $\textbf{30}\times$ less context length.
@@ 5.3 Results on Demonstration Selection
Finally, compared with top-$k$ and random-$k$ for $k$ up to $150$, we find that our approach matches top-$k$ when $k=150$ ($9000$ tokens) with $10$ in-context examples ($600$ tokens). See illustration in Figure 4(b). In this setting, our approach requires only $13$ GB of GPU memory, whereas top-$k$ and random-$k$ use $47$ GB to achieve similar performance, all tested on DeepSeek-7B. Our method also performs well on other models, such as Llama-8B and Qwen-7B. On the three reasoning tasks, ICL-GradRE outperforms the most competitive baseline by $22\%$ on average, while using $49\%$ less computation.
@@ 6 Related Work
Building on these insights, a growing line of work has focused on improving demonstration selection and usage. ADAPT Ross and Andreas (2024) introduces a framework in which a teacher model diagnoses student misconceptions and adaptively selects demonstrations. Treating demonstration selection as a retrieval problem, Yu et al. (2025) propose a reinforcement learning–based ranking method that updates retrieval rankings using feedback from the LLM, enabling better performance on long-tail samples. Wan et al. (2025) take a different perspective by fitting a Gaussian process regression model that maps subsets of demonstrations to model losses. Another direction involves ensemble-based strategies that combine multiple subsets of demonstrations. Khalifa et al. (2023) conduct inference separately on multiple subsets and aggregate the predictions using weighted ensembles determined by input similarity. Similarly, Huang et al. (2024) partition demonstrations into groups and reweight output logits through a non-gradient optimization procedure.
It is also worth noting earlier research on feature selection in unsupervised learning, such as methods based on expectation–maximization and clustering Dy and Brodley (2004). Recent work has advanced the theoretical understanding of in-context learning (ICL). Zhang et al. (2024b) analyze transformers with a single linear self-attention layer and show that such models can in-context learn linear functions. Wu et al. (2024) establish complexity bounds for attaining Bayes-optimal ICL with single-layer linear attention models. We refer readers to the recent survey by Dong et al. (2022) for a more comprehensive review of in-context learning.

### v2

@@ Abstract
This paper introduces an algorithm to select demonstration examples for in-context learning of a query set. Given a set of $n$ examples, how can we quickly select $k$ out of $n$ to best serve as the conditioning for downstream inference? This problem has broad applications in prompt tuning and chain-of-thought reasoning. Since model weights remain fixed during in-context learning, previous work has sought to design methods based on the similarity of token embeddings. This work proposes a new approach based on gradients of the output taken in the input embedding space. Our approach estimates model outputs through a first-order approximation using the gradients. Then, we apply this estimation to multiple randomly sampled subsets. Finally, we aggregate the sampled subset outcomes to form an influence score for each demonstration, and select $k$ most relevant examples. This procedure only requires pre-computing model outputs and gradients once, resulting in a linear-time algorithm relative to model and training set sizes. Extensive experiments across various models and datasets validate the efficiency of our approach. We show that the gradient estimation procedure yields approximations of full inference with less than ${1}\%$ error across six datasets. This allows us to scale up subset selection that would otherwise run full inference by up to ${37.7}\times$ on models with up to $34$ billion parameters, and outperform existing selection methods based on input embeddings by ${11}\%$ on average.
@@
footnotetext: *Equal Contribution. Email correspondence can be directed to all authors at {zhang.zini, zhang.zhens, li.dongyu, ho.zhang}@northeastern.edu, wangluxy@umich.edu, and jdy@ece.neu.edu.
@@ 1 Introduction
Table 1: Runtime comparison between our algorithms and standard subset selection methods for demonstration selection. Here, $T$ represents the time cost of running one forward pass on a model with the training set, $k$ denotes the number of selected examples, and $n$ represents the number of demonstrations. Our approach computes model outputs with one forward pass of time $T$ and input gradients with time $2T$, plus an overhead cost of $O(n\log n)$, which is negligible relative to $T$.
@@ 2 Preliminaries
Let $p$ denote the embedding dimension. Our goal is to select a subset $S\subseteq\mathcal{S}^{\text{Demo}}$ that minimizes the above objective. After selecting a set of demonstrations, we evaluate the test performance on a separate test set.
@@ 3 Scaling Up Inference via Estimation
A case study of linear functions. We illustrate the above approximation with an example of learning linear functions with transformer models Garg et al. (2022). We use this setting to evaluate the accuracy of equation (1) for ICL of linear functions. Suppose that both the in-context example and query example are generated as $y_{i}=\left\langle\beta^{(j)},x_{i}\right\rangle+\epsilon_{i}$, where the input vector $x_{i}$ is randomly sampled from $\mathcal{N}(0,\operatorname{Id}_{p})$ with $p=20$. The coefficient vector $\beta^{(j)}$ is randomly drawn from a finite set $\{\beta^{(1)},\beta^{(2)},\cdots,\beta^{(C)}\}$, meaning that different in-context examples may have different coefficients. We consider two settings for the noise term $\epsilon_{i}$: linear regression with $\epsilon_{i}=0$ for all $i$, or noisy linear regression with $\epsilon_{i}\sim\mathcal{N}(0,1)$.
Figure 2(a) shows that in both settings, ICL-GradEx can accurately estimate the true outcomes of ICL. In particular, as $k$ increases, the trained transformer learns the linear function correctly, and the residual sum of squares (RSS) between the estimated scores and the true scores gradually decreases to near zero. Recall that $p=20$. We thus plot from $k=p$ onward because this is the minimum number of samples required to solve a system of linear equations. When $k<p$, the trained transformer is unable to fully determine the linear function. (See Figure 2, Garg et al. (2022)).
@@ 4 Efficient Demonstration Selection
Figure 2: An illustration of our approach for in-context learning of linear functions, as we vary the number of in-context examples $k$. Figure 2(a): ICL-GradEx incurs low approximation error relative to full inference, for both linear and nonlinear regression. Figure 2(b): With ICL-GradSel, the selected demonstrations follow the same linear function $\beta$, achieving lower error than top-$k$ and random-$k$ selections.
@@ 5 Experiments
Our experiments show that, across a wide range of LLMs (such as Llama, OPT, DeepSeek, and Qwen), ICL-GradEx can estimate model inference outcomes within $\textbf{1}\%$ error, for LLMs ranging from $1$ billion up to $34$ billion parameters. Crucially, this estimation allows us to reduce the computation cost of subset selection by up to $\textbf{37.7}\times$. When applied to in-context learning across six datasets that include sentence classification and math reasoning tasks, ICL-GradRE improves the test $F_{1}$ score by $\textbf{11}\%$ on average. For long-context settings where $k=150$, our selection matches the performance of top-$k$ and random-$k$ selection with $\textbf{30}\times$ less context length.
@@ 5.3 Results on Demonstration Selection
Finally, compared with top-$k$ and random-$k$ for $k$ up to $150$, we find that our approach matches top-$k$ selection when $k=150$ ($9000$ tokens) with $10$ in-context examples ($600$ tokens). See illustration in Figure 4(b). In this setting, our approach requires only $13$ GB of GPU memory, whereas top-$k$ and random-$k$ selection use $47$ GB to achieve similar performance, all of which are tested on the DeepSeek-7B model. Our method also performs well on other models, such as Llama-8B and Qwen-7B. On the three reasoning tasks, ICL-GradRE outperforms the most competitive baseline by $22\%$ on average, while using $49\%$ less computation.
@@ 6 Related Work
Building on these insights, a growing line of work has focused on improving demonstration selection and usage. ADAPT Ross and Andreas (2024) introduces a framework in which a teacher model diagnoses student misconceptions and adaptively selects demonstrations. Yu et al. (2025) propose a reinforcement learning–based ranking method that updates retrieval rankings using feedback from the LLM, enabling better performance on long-tail samples. Wan et al. (2025) take a different perspective by fitting a Gaussian process regression model that maps subsets of demonstrations to model losses. Another direction involves ensemble-based strategies that combine multiple subsets of demonstrations. Khalifa et al. (2023) conduct inference separately on multiple subsets and aggregate the predictions using weighted ensembles determined by input similarity. Similarly, Huang et al. (2024) partition demonstrations into groups and reweight output logits through a non-gradient optimization procedure.
It is also worth noting earlier research on feature selection in unsupervised learning, such as methods based on expectation-maximization and clustering Dy and Brodley (2004). Recent work has advanced the theoretical understanding of in-context learning (ICL). Zhang et al. (2024b) analyze transformers with a single linear self-attention layer and show that such models can in-context learn linear functions. Wu et al. (2024) establish complexity bounds for attaining Bayes-optimal ICL with single-layer linear attention models. We refer readers to the recent survey by Dong et al. (2022) for a more comprehensive review of in-context learning.